# Professionally translated parallel data for 115 under-represented languages

### v1

@@ 5.1 Evaluation
Since so many language pairs are covered, we evaluate on a combination of all available evaluation sets, namely Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019), and Gatones (Jones et al., 2023). Since no reliable embedding models exist for these languages, we use ChrF with NFKC unicode normalization as our metric.
N-shot: ChrF-counterweighted RAG
To have a strong baseline for N-shot results, we adopt a RAG-based approach that resembles the greedy set-cover algorithm to choose the few-shot exemplars. For each sentence in the eval set, we want the best coverage of the source sentence n-grams as possible, with the least redundancy among exemplars. Therefore, we iteratively choose the exemplar whose source side has the minimum ChrF to the eval source. However, when counting the true positives in the ChrF calculation, we weight the count of each ngram by , where is the number of times has been seen among the exemplars chosen so far, and is a parameter to control how close this algorithm is to ngram set-cover. We use . The set of exemplars we choose from is the concatenation of SmolSent and SmolDocSplit.
@@ 5.2 Finetuning Setup and Results
We finetuned Gemini 2.0 Flash for 50 epochs on SmolDoc, SmolSent, a combination of the two (Both), and their combination plus Gatitos (Both+G). To simplify finetuning, we split SmolDoc into sentence pairs (SmolDocSplit). We use a minimalist prompt (Appendix C).
Results can be seen in Table 3. Finetuning on SmolSent gives an average gain of +2.7 ChrF, and SmolDocSplit give the exact same gain on its languages. Mixing the two leads to a gain of +3.3 to +3.6 ChrF, and adding in Gatitos bumps it up to +3.6 to +4.1 ChrF, passing all other 0-shot baselines except Google Translate (on the languages it supports).
The 10-shot RAG results on the un-tuned model are very close to the finetuned 0-shot results, and the finetuned models show no benefit from multi-shot decoding, suggesting that these are two different ways of giving the same information—inference-time versus training time. The 10-shot random results (not included in table) were much lower.
@@ 5.3 The Problem with xxen training
Our initial experiments used all data for both enxx and xxen . However, the models lost performance on all tasks. The root cause turned out to be the multiway-parallel data with English targets. Large language models are especially susceptible to repetition in data (Lee et al., 2022), and with 115 language pairs, for every one epoch over the data, the model saw about 115 epochs for each individual target sentence. Therefore, it wildly overfit and lost performance on all language pairs. Mitigating such overfitting is an important research direction to pursue, since many promising datasets are multiway parallel, e.g. Flores-101 (Goyal et al., 2022), Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019), and others. However, this is out of scope for the present paper, so we restrict our experiments to enxx .
@@ 6 Factuality Review
Since SmolDoc contains LLM-generated sources, they contain some factual inaccuracies. We therefore do a full human audit and assign factuality codes to each document. Each of the 584 documents was rated by three raters. Each rating is accompanied by a detailed explanation, including sources cited. Inter-annotator agreement was high, with Cohen’s between each pair of raters between 0.82 and 0.87. The error code distribution can be seen in Fig. 2. Appendix D gives details on the rubric.
@@ 7 Conclusion
We have open-sourced the Smol dataset, a professionally-translated dataset covering 115 low resource languages and targeting the tasks of translation and factuality. It comprises SmolDoc and SmolSent, two training datasets with the complementary strengths of sentence selection (complex, and high token coverage) and document generation (contextual, varied domains, simpler sentences) respectively. We demonstrate that finetuning Gemini 2.0 Flash on these yields to substantial improvements in translation quality. Smol joins a growing body of resources to support underserved languages in the age of AI.
@@ 8 Limitations
9 Ethical implications and Potential Risks
Our released Smol dataset will assist in development and evaluation of machine translation and multilingual AI capabilities for low-resource languages. Although we consider this a net benefit for society, one potential risk is that users may overestimate the quality of machine translations in these languages. Another potential risk is increasing the proliferation of AI-generated content in the target languages on the open web, which may pose issues for future development of web-crawled datasets in these languages. Furthermore, although we have surveyed and talked with native speakers of all Smol languages and found only enthusiasm for technological support of their languages, communities are not monoliths and some community members may prefer that technology not support their language.
Appendix A SmolSent details
A.1 Evaluating the SmolSent selection process
In Section 3.1.1, we describe experiments used to validate the selection process for SmolSent. We train a backbone MT system is pretrained on the Madlad-400 dataset (Kudugunta et al., 2023). The following languages are held out of the training data to be used for fine-tuning experiments: Catalan, Icelandic, Marathi, Turkish, Maltese, Xhosa, Tamil, Basque, and Tajik. The model itself is a 1B parameter encoder-decoder Transformer and is trained from scratch on the Madlad-400 data. Each of the candidate data selection methods is used to select data from the held-out languages, and then each candidate set is used in turn to finetune the backbone model. The results of this finetune step are reported in Table 1, where the set-cover approach is shown to be most effective.
A.2 Notes on Researcher in the Loop
@@
As described in the paper text, this approach is designed to combat issues such as honeypot sentences. Example Honeypot sentences can be seen in Table 4
@@
A.3 More detailed corpus statistics on SmolSent
To measure “Bang for our Buck” we define the excess-token ratio as the number distinct tokens in the set cover divided by the number of target tokens, and use it along with the coverage percent to understand the SmolSent dataset. Table 5 compares corpus statistics of SmolSent to four other corpora. sametoks picks a random set of sentences from CommonCrawl until it has the same number of tokens as SmolSent; this only covers 50% of the target tokens and has an excess-token-ratio ratio of 3.3, much worse than SmolSent’s value of 2.3. The samecov baseline randomly picks common-crawl sentences until it has the same token coverage as SmolSent, which necessitates a 67x larger set of sentences, and a correspondingly bloated excess token coverage ratio of 23.1. As a further reference we compare tiers 1 (largest) and 5 (smallest; comparative size to SmolSent) of SmolDoc. As expected of machine-generated text, they have a worse value, corresponding to a narrower spectrum of vocabulary used.
@@ A.3 More detailed corpus statistics on SmolSent
Appendix B SmolDocDetails
B.1 SmolDoc data tier details
Per-tier statistics on the SmolDoc dataset can be seen in Table 6.
@@ B.1 SmolDoc data tier details
SmolDoc-t5 | 34 | 2.2k | 401k | 2.2M | 66 | 12k | 65k |
all-Smol | 115 | 97k | 6.1M | 34M | 827 | 52k | 294k |
B.2 Details on SmolDoc prompt creation
@@
one of 600 manually created topics, e.g. "volcanic eruptions" or "A special tree"
one of 50 tone/tense categories, e.g. "Please use the subjunctive mood.", "Use an effusive tone.""
A style prompt, e.g. "You are the author R.K. Narayan." or "You are a mother talking to her son."
@@
For each prompt, we generated 8 responses (T=0.7). These were ranked by their simple token density (unique tokens over total tokens), and the top two were chosen for consideration. Using the Researcher-in-the-loop mentality (“measure twice cut once!"), we went over 1000+ responses by hand and scored/edited them. This was mainly to filter out questionable or boring responses. A typical paragraph scored as 0 would be LLM-speak like “X is a complex and multifaceted problem with no easy solution. Here are some suggestions. Keep in mind that there is no one-size-fits all solution, and ultimately, the choice is up to you. […]”.
Example prompts can be seen in Table 7.
@@
You are a teenager talking to his friend. Please carefully craft a 1-paragraph bit about an engineer who subsists off coffee. Try to include the words "confirmed", "move" and "above". |
@@
Please write a long passage starting with ’Mum and Dad pause their debate when we hear this creepy clacking that sounds like hail falling.’ |
@@ B.3 SmolDoc Errata
Appendix C Prompts for Decoding
@@
You are an expert translator. I am going to give you some example pairs of text snippets where the first is in ${SL} and the second is a translation of the first snippet into ${TL}. The sentences will be written ${SL}: <first sentence> ${TL}: <translated first sentence> After the example pairs, I am going to provide another sentence in ${SL} and I want you to translate it into ${TL}. Give only the translation, and no extra commentary, formatting, or chattiness. Translate the text from ${SL} to ${TL}.
@@
Appendix D Factuality rubric
The full factuality rubric given to raters is presented in Table 8.
Rating | Definition of Rating |
N/A - Not Applicable | True/False does not apply here. Most stories, dialogues, or fictional works would be considered N/A, unless they are promoting a falsehood about the real world. |
No Issues | All claims are factual and accurate. (It’s OK if they are out-of-date, e.g. “Barack Obama is the President of the USA”) |
Minor Issue(s) | There are small inaccuracies. For instance, it may be broadly correct but frame something in a misleading way. |
Clear Issue(s) | There are clear mistakes in factuality. |
Appendix E Translations for NKo
@@
Appendix F Full results
@@
lang | cat | G2F | +sS | +sD | +sB | +sG | Cld | +RAG | G2P | GPT4o | GTr | NLLB | |
ee | Both | +36.1 | 3.0 | 37.7 | 37.5 | 39.1 | 39.2 | 37.8 | 40.0 | 39.5 | 7.5 | 42.7 | 40.7 |
kr | Both | +10.8 | 17.3 | 25.6 | 26.3 | 28.1 | 28.8 | 22.7 | 26.3 | 20.3 | 22.2 | 32.6 | 31.0 |
kg | Both | +9.2 | 34.9 | 46.9 | 35.6 | 44.1 | 43.2 | 43.2 | 47.0 | 37.8 | 29.1 | 50.2 | 3.4 |
bem | Both | +7.3 | 40.0 | 44.8 | 45.1 | 47.3 | 49.2 | 43.3 | 47.7 | 42.3 | 33.3 | 49.7 | 41.8 |
dyu | Both | +5.3 | 17.9 | 22.5 | 22.9 | 23.2 | 23.7 | 23.9 | 24.4 | 21.0 | 4.5 | 22.4 | 12.5 |
din | Both | +4.6 | 20.3 | 23.8 | 23.1 | 24.9 | 25.1 | 23.3 | 25.9 | 21.4 | 1.6 | 25.1 | 26.5 |
st | Both | +4.5 | 49.9 | 54.0 | 55.3 | 54.4 | 55.0 | 49.4 | 57.0 | 53.1 | 49.2 | 49.0 | 47.2 |
luo | Both | +4.1 | 37.4 | 39.1 | 41.3 | 41.5 | 42.0 | 39.1 | 42.0 | 39.6 | 36.1 | 41.3 | 39.5 |
fon | Both | +3.6 | 21.3 | 24.3 | 24.1 | 24.9 | 25.3 | 20.4 | 23.7 | 23.8 | 1.9 | 25.9 | 24.2 |
bm | Both | +3.4 | 30.8 | 28.6 | 34.9 | 34.2 | 34.1 | 34.0 | 36.2 | 33.9 | 9.0 | 35.7 | 32.2 |
ak | Both | +2.6 | 35.5 | 36.1 | 37.6 | 38.1 | 38.2 | 34.4 | 38.1 | 37.3 | 32.2 | 34.5 | 33.3 |
ln | Both | +2.5 | 46.8 | 48.1 | 49.2 | 49.3 | 49.3 | 44.6 | 48.3 | 46.5 | 45.2 | 46.4 | 45.7 |
wo | Both | +1.1 | 30.3 | 30.0 | 30.8 | 31.4 | 31.6 | 31.4 | 32.2 | 30.7 | 29.8 | 36.2 | 30.9 |
tum | Both | +1.1 | 40.8 | 39.5 | 41.9 | 41.9 | 42.8 | 40.0 | 42.7 | 43.8 | 37.7 | 45.4 | 36.2 |
ff | Both | +0.9 | 25.0 | 24.4 | 25.7 | 25.9 | 26.5 | 25.7 | 26.1 | 25.2 | 2.5 | 25.9 | 27.1 |
ti | Both | +0.8 | 24.2 | 24.3 | 24.6 | 25.0 | 25.7 | 25.0 | 26.2 | 26.1 | 9.3 | 26.1 | 25.5 |
yo | Both | +0.6 | 34.6 | 33.8 | 35.4 | 35.2 | 35.8 | 29.2 | 36.8 | 26.7 | 27.6 | 21.3 | 32.6 |
tn | Both | +0.2 | 52.5 | 50.8 | 51.1 | 52.7 | 53.2 | 50.1 | 51.7 | 53.3 | 36.8 | 55.6 | 53.0 |
am | Both | +0.1 | 34.0 | 33.2 | 33.1 | 34.1 | 33.5 | 31.6 | 32.6 | 35.8 | 29.6 | 34.7 | 30.3 |
ig | Both | -0.1 | 47.2 | 47.1 | 47.7 | 47.1 | 47.8 | 43.9 | 46.2 | 47.8 | 46.2 | 47.6 | 46.6 |
so | Both | -0.1 | 49.7 | 46.2 | 50.1 | 49.6 | 49.1 | 48.7 | 49.8 | 50.3 | 50.8 | 50.6 | 48.6 |
sn | Both | -0.5 | 50.5 | 48.3 | 49.9 | 50.0 | 50.3 | 46.8 | 48.8 | 51.8 | 50.3 | 49.2 | 48.2 |
ss | Both | -0.6 | 50.2 | 48.8 | 47.7 | 49.6 | 50.3 | 49.6 | 51.2 | 51.8 | 46.0 | 56.3 | 48.1 |
yue | Both | -0.7 | 26.8 | 25.8 | 25.6 | 26.1 | 26.1 | 28.2 | 28.2 | 27.5 | 31.6 | 25.9 | 22.6 |
om | Both | -0.8 | 40.1 | 38.0 | 39.6 | 39.3 | 39.4 | 39.0 | 40.2 | 41.3 | 38.4 | 41.4 | 39.1 |
lg | Both | -1.1 | 42.5 | 39.9 | 41.0 | 41.4 | 41.7 | 42.0 | 43.1 | 43.5 | 41.0 | 43.6 | 41.1 |
rw | Both | -1.4 | 45.2 | 43.1 | 43.1 | 43.8 | 43.8 | 43.2 | 44.0 | 45.1 | 44.7 | 48.8 | 43.4 |
sw | Both | -1.5 | 66.7 | 64.6 | 64.7 | 65.2 | 64.8 | 64.0 | 65.5 | 67.2 | 66.5 | 65.3 | 60.5 |
mg | Both | -1.9 | 52.8 | 48.9 | 51.1 | 50.9 | 51.5 | 52.4 | 52.5 | 53.3 | 52.2 | 52.6 | 52.1 |
zu | Both | -2.0 | 58.3 | 56.4 | 55.2 | 56.3 | 56.1 | 54.1 | 55.5 | 58.5 | 57.5 | 57.6 | 57.6 |
xh | Both | -2.3 | 53.9 | 49.7 | 51.4 | 51.6 | 51.8 | 51.6 | 52.3 | 53.9 | 53.7 | 54.8 | 51.2 |
nso | Both | -2.3 | 46.8 | 42.8 | 44.0 | 44.5 | 44.6 | 46.9 | 47.7 | 48.1 | 46.9 | 47.6 | 45.5 |
ber | Both | -3.2 | 25.3 | 20.6 | 23.0 | 22.1 | 21.9 | 28.5 | 25.2 | 31.1 | 2.8 | 21.0 | 32.4 |
ha | Both | -3.4 | 54.5 | 50.7 | 50.5 | 51.1 | 51.5 | 50.9 | 51.0 | 54.1 | 53.9 | 53.8 | 53.9 |
ts | Both | -3.6 | 50.6 | 47.2 | 46.5 | 47.0 | 48.1 | 49.7 | 50.1 | 51.6 | 49.0 | 52.9 | 51.3 |
rn | Both | -3.9 | 44.5 | 39.3 | 40.7 | 40.6 | 40.9 | 43.1 | 43.4 | 46.2 | 44.3 | 45.4 | 45.0 |
af | Both | -4.2 | 71.9 | 68.8 | 67.0 | 67.7 | 68.3 | 71.7 | 72.5 | 72.1 | 71.8 | 71.5 | 68.6 |
ny | Both | -5.1 | 55.0 | 47.5 | 50.2 | 49.9 | 49.7 | 53.0 | 53.1 | 55.3 | 53.9 | 55.8 | 50.3 |
[…]

### v2

@@ 5.1 Evaluation
Since so many language pairs are covered, we evaluate on a combination of all available evaluation sets, namely Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019),
and an in-house eval set. Since no reliable embedding models exist for these languages, trained metrics are not an option, so we use ChrF Popović (2015) as implemented in SacreBleu (Post, 2018)111111signature: case.mixed+numchars.6+numrefs.1
+space.False+tok.none+version.1.3.0 with NFKC unicode normalization as our metric. For ten-shot decoding, exemplars were selected from both sub-datasets of Smol using ChrF-counterweighted RAG (Appendix D).
@@ 5.2 Finetuning Setup and Results
We finetuned Gemini 2.0 Flash for 40 epochs on SmolDoc, SmolSent, a combination of the two (Both), and their combination plus Gatitos (Both+G). To simplify finetuning, we split SmolDoc into sentence pairs (SmolDocSplit).
Results can be seen in Table 3. Finetuning on SmolSent gives an average gain of +2.7 ChrF points, and SmolDocSplit gives +2.6 ChrF points on its languages. Concatenating the two training datasets leads to a gain of +3.3 to +3.6 ChrF points, and adding in Gatitos bumps it up to +3.8 to +4.1 ChrF points, passing all baselines except Google Translate. The 10-shot RAG results on the un-tuned model are very close to the finetuned 0-shot results, and the finetuned models show no benefit from multi-shot decoding, suggesting that these are two different ways of giving the same information—inference-time versus training time. The 10-shot random results (not included in table) were much lower.
Gains were highest on languages that are not related to mid- or high-resource languages, and lowest on dialects close to major languages. As a heuristic to measure this, we exclude languages that are on Google Translate or closely related to languages on it (Appendix G). The average gain on these languages jumps to +7.6 ChrF.
@@ 5.3 The Problem with xxen training
Our initial experiments used all data for both enxx and xxen . However, the models lost performance on all tasks. The root cause turned out to be the multiway-parallel data with English targets. LLMs are especially susceptible to repetition in data (Lee et al., 2022), and with 115 language pairs, for every one epoch over the data, the model saw about 115 epochs for each individual target sentence. Therefore, it wildly overfit and lost performance on all language pairs. Mitigating such overfitting is an important research direction to pursue, since many promising datasets are multiway parallel, e.g. Flores-101 (Goyal et al., 2022), Flores-200 (Team et al., 2022), NTreX Federmann et al. (2022b); Barrault et al. (2019), and others. However, this is out of scope for the present paper, so we restrict our experiments to enxx .
@@ 6 Factuality Review
Since SmolDoc contains LLM-generated sources, they contain some factual inaccuracies. We therefore do a full human audit and assign factuality codes to each document. Each of the 584 documents was rated by three raters. Each rating is accompanied by a detailed explanation, including sources cited. Inter-annotator agreement was high, with Cohen’s between each pair of raters between 0.82-0.87. The error code distribution can be seen in Figure 2. The rubric is presented in Table 4.
| Rating | Definition of Rating |
| N/A | True/False does not apply here. Most stories, dialogues, or fictional works would be considered N/A, unless they are promoting a falsehood about the real world. |
| No Issues | All claims are factual and accurate. (Out-of-date is fine, e.g. “Barack Obama is the US President”) |
| Minor Issue(s) | There are small inaccuracies. E.g., it may be broadly correct but frame something in a misleading way. |
| Clear Issue(s) | There are clear mistakes in factuality. |
@@ 7 Conclusion
We have open-sourced the Smol dataset, a professionally-translated dataset covering 124 (and growing!) low resource languages and targeting the tasks of translation and factuality. It comprises SmolDoc and SmolSent, two training datasets with the complementary strengths of sentence selection (complex, and high token coverage) and document generation (contextual, varied domains, simpler sentences) respectively. We demonstrate that finetuning Gemini 2.0 Flash on these yields to substantial improvements in translation quality. Smol joins a growing body of resources to support underserved languages in the age of AI.
@@ 8 Limitations
Appendix A Operational Definition of LRL
In this paper, we operationally define LRL as any language beyond the first 100 supported by most traditional crawls and MT providers. Since there is some variation in which languages exactly this is, we concretize it as the set of 104 languages supported in Google Translate prior to 2020. These are the languages for which launchable quality was possible before LLM-type models like M4 (Arivazhagan et al., 2019; Bapna et al., 2022) and PaLM (Anil et al., 2023) came on these scene. It is also worth noting that since these languages were on the product for much longer, they have much more machine-translated content online from services that used Google Translate for internationalization. These 104 languages are: af am ar az be bg bn bs ca ceb co cs cy da de el en eo es et eu fa fi fil fr fy ga gd gl gu ha haw hi hmn hr ht hu hy id ig is it he ja jv ka kk km kn ko ku ky la lb lo lt lv mg mi mk ml mn mr ms mt my ne nl no ny pa pl ps pt ro ru sd si sk sl sm sn so sq sr st su sv sw ta te tg th tr uk ur uz vi xh yi yo zh zh-Hant zu.
Rightly speaking, the languages outside of this set might better be termed "Very Low-Resource" instead of just "Low Resource", since the 104 languages above do include languages like Hawaiian, Javanese, Yiddish, and Hmong, which are by no stretch of the imagination high-resource. We will leave more rigorous definitions to future work.
Appendix B SmolSent details
B.1 Evaluating the SmolSent selection process
In Section 3.1, we describe experiments used to validate the selection process for SmolSent. We train a backbone MT system is pretrained on the Madlad-400 dataset (Kudugunta et al., 2023). The following languages are held out of the training data to be used for fine-tuning experiments: Catalan, Icelandic, Marathi, Turkish, Maltese, Xhosa, Tamil, Basque, and Tajik. The model itself is a 1B parameter encoder-decoder Transformer and is trained from scratch on the Madlad-400 data. Each of the candidate data selection methods is used to select data from the held-out languages, and then each candidate set is used in turn to finetune the backbone model. The results of this finetune step are reported in Table 1, where the set-cover approach is shown to be most effective.
B.2 Notes on Researcher in the Loop
@@
As described in the paper text, this approach is designed to combat issues such as honeypot sentences. Example Honeypot sentences can be seen in Table B.1
@@
B.3 Corpus statistics on SmolSent
To measure “Bang for our Buck” we define the excess-token ratio as the number distinct tokens in the set cover divided by the number of target tokens, and use it along with the coverage percent to understand the SmolSent dataset. Table B.2 compares corpus statistics of SmolSent to four other corpora. sametoks picks a random set of sentences from CommonCrawl until it has the same number of tokens as SmolSent; this only covers 50% of the target tokens and has an excess-token-ratio ratio of 3.3, much worse than SmolSent’s value of 2.3. The samecov baseline randomly picks common-crawl sentences until it has the same token coverage as SmolSent, which necessitates a 67x larger set of sentences, and a correspondingly bloated excess token coverage ratio of 23.1. As a further reference we compare tiers 1 (largest) and 5 (smallest; comparative size to SmolSent) of SmolDoc. As expected of machine-generated text, they have a worse value, corresponding to a narrower spectrum of vocabulary used.
@@ B.3 Corpus statistics on SmolSent
Appendix C SmolDocDetails
C.1 SmolDoc data tier details
Per-tier statistics on the SmolDoc dataset can be seen in Table C.1.
@@ C.1 SmolDoc data tier details
| SmolDoc-t5 | 38 | 2.5k | 448k | 2.5M | 66 | 12k | 65k |
C.2 Details on SmolDoc prompt creation
@@
one of 600 manually created topics, e.g. “volcanic eruptions” or "A special tree”
one of 50 tone/tense categories, e.g. “Please use the subjunctive mood.”, “Use an effusive tone.”
A style prompt, e.g. “You are the author R.K. Narayan.” or “You are a mother talking to her son.”
@@
For each prompt, we generated 8 responses (T=0.7). These were ranked by their simple token density (unique tokens over total tokens), and the top two were chosen for consideration. Using the Researcher-in-the-loop mentality (“measure twice cut once!”), we went over 1000+ responses by hand and scored/edited them. This was mainly to filter out questionable or boring responses. A typical paragraph scored as 0 would be LLM-speak like “X is a complex and multifaceted problem with no easy solution. Here are some suggestions. Keep in mind that there is no one-size-fits all solution, and ultimately, the choice is up to you. […]”.
Example prompts can be seen in Table C.2.
@@
| You are a teenager talking to his friend. Please carefully craft a 1-paragraph bit about an engineer who subsists off coffee. Try to include the words “confirmed”, “move” and “above”. |
@@
| Please write a long passage starting with ‘Mum and Dad pause their debate when we hear this creepy clacking that sounds like hail falling.’ |
@@ C.3 SmolDoc Errata
Appendix D N-shot: ChrF-counterweighted RAG
To have a strong baseline for N-shot results, we adopt a RAG-based approach that resembles the greedy set-cover algorithm. For each sentence in the eval set, we want the best coverage of the source sentence -grams as possible, with the least redundancy among exemplars. Therefore, we iteratively choose the exemplar whose source side has the minimum ChrF to the eval source. However, when counting the true positives in the ChrF calculation, we weight the count of each ngram by , where is the number of times has been seen among the exemplars chosen so far, and is a parameter to control how close this algorithm is to ngram set-cover. We use .The set of exemplars we choose from is the concatenation of SmolSent and SmolDocSplit.
Appendix E Prompts for Decoding
@@
You are an expert translator. I am
going to give you some example
pairs of text snippets where the
first is in ${SL} and
the second is a translation of
the first snippet into
${TL}. The sentences
will be written
${SL}: <first sentence>
${TL}: <translated first sentence>
After the example pairs, I am
going to provide another sentence
in ${SL} and I want
you to translate it into
${TL}. Give only the
translation, and no extra
commentary, formatting, or
chattiness. Translate the text
from ${SL} to ${TL}.
@@
Appendix F Volunteer contributions
A few languages have extra details that need to be called out here.
F.1 Translations for Cantonese
A volunteer team of Cantonese speakers at Google pulled together to translate the maximal set of Smol text. Mingfei Lau and Jonathan Eng were the main leaders of this effort, and the contributors to translation and post-editing were (alphabetically): Tsz Yan Au, Emily Awesome, Jason Chan, Siu Man Chan, Vicky Chan, Yiwang Chen, Kinton Cheung, Mingo Choi, Andy Chow, Ashley Chow, Olivia Chow, Daniel (Ying Wai) Fan, Thomas Fung, Vikki Ha, Joshua Kwong, Liam Lee Pong Lam, Jonas Lau, Ying Tung (Grace) Law, Crystal Lee, Aki Leung, Derek Leung, Jackie Leung, Thomas Leung, Mu Li, Alicia Liu, Malena Loosli, Chui McConnell, Ken Ng, Nicholas Ng, Tonia Shen, Helen Shum, Franky Sze, Eric Tang, Tommy Tse, Daniel Wong, Danny Wong, Maggie Wong, Pinki Wong, Jeffrey Yu, Shanelle Yu, Shing Fung Yue, Miranda Zhang, and Willis Zhang.
F.2 Translations for NKo
@@
F.3 Translations for Zazaki, Hawrami, and Gilaki
Sina Ahmadi gratefully acknowledges support from the UZH Postdoc Grant (reference number 269093).
F.4 Translations for Zarma
Annotation Pipeline
The Zarma translation process of SMOL—all the subsets—was done through a combination of automatic and human in the loop methods. We leveraged some existing tools that our team developed to speed up the annotation process. We first used a baseline bidirectional model that we developed to produce initial translation of the samples. These machine translated samples were then passed through our Zarma grammatical error correction model. This model was built by pre-training gemma-2-9b on Zarma data and fine tuning the checkpoint on grammar error correction data set using Direct Preference Optimization (DPO) settings. The outputs from this stage—both languages side by side—were then given to our team of annotators for review.
The annotators were given some guidelines—in addition to the general guidelines from Smol—for the annotations. These guidelines include:
Word adaptation: rules for handling technical terms, proper nouns, and domain-specific vocabulary that might not have direct equivalents in Zarma. E.g: all the scientific/technical words remain unchanged; and words that have known french-ized equivalent in Zarma must be used in their french-ized forms (for better understandability).
Prioritize understandability: guidelines to prioritize understandability and fidelity over word-for-word translation. We instructed annotators to focus on creating translations that sound natural and widely understandable by Zarma speakers.
Language specific constraints: language specific guidelines that cannot be generalized.
The pipeline speeds up the process while maintaining the quality, since some of the outputs from the automatic stages were already correct.
Zarma Community Attitude Towards Tech
: The Zarma community—and the whole Niger in general—are very open minded regarding technology. When we started our very first resource creation for the Zarma language, we received positive feedback and even help from the community, as long as we developed an openly accessible solution for the community. For the SMOL annotation, that trust helps us to receive valuable help. For instance, a government based institution verbally promised to accompany any language preservation—machine learning focus in our case—if the outcome will be open-sourced for community usage.
F.5 Post-Edits for Mooré
Annotation Pipeline
The annotation process for Mooré did not involve any automated components; everything was annotated by humans. The annotation focuses more on the guidelines provided by Smol, in addition to some more as in the Zarma case.
Mooré Community Perspective
The Mooré community, similarly to the zarma community, are very open minded towards technology; especially if it touches cultural/language preservation. One main feedback we received from some elders (parents of one of the annotators) was a warning to ONLY USE standard Mooré orthography, not any equivalent. They want the language to be well documented according to the language standards.
F.6 Post-Edits for Indonesian
A volunteer team post-edited translations of smoldoc and smolsent datasets that had done by Gemini 2.5 Pro. The contributors to translation and post-editing were Muhammad Ravi Shulthan Habibi, David Anugraha, and Genta Indra Winata. The post-edits resulted in about 70% of the machine translations being changed.
Translators agreed that the system output was often too “formal", “stiff", or “awkward". The “formal" translations were furthermore not formal in an acceptible sense, but “too awkward and stiff, even for a more formal situations", as an annotator said. Each word choice might be correct and standard in Indonesian, but when combined in a sentence, the result sounded unnatural. Therefore, the majority of the post-edits focused on making the translations sound more natural.
Nonetheless, overall the system output was already quite reasonable in terms of register. In some cases, though, it leaned toward being too rigid. The post-edits tried to loosen that into a consistent “medium” range, but with some flexibility depending on the style of each sentence (sometimes slightly more formal, sometimes slightly less) so the overall text still feels natural and coherent.
F.7 Translations for Languages of the Russian Federation
Traditionally, speakers of hundreds of Cyrillic-based languages in the Russian Federation translate datasets via Russian. For the success of this project, I (Ali Kuzhuget) first funded a professional translation into Russian, engaging Andrey Anisimov as the main translator. The proofreading was conducted by Farhad Fatkullin, Vice-President of the National League of Translators, together with machine translation specialist David Dalé. I also oversaw formatting correctness and coordinated the overall translation workflow.
In parallel, I supervise the translation of the dataset from Russian into Tuvan, using a dedicated Telegram chatbot for large-scale dataset translation. This tool enables multiple rounds of validation and systematic assessment of translation quality. Currently, representatives of about a dozen Cyrillic languages are in the process of translating the Smol dataset into their own languages through Russian and/or English (for example, Tuvan, Bashkir, Chuvash, and others), ensuring both linguistic accuracy and cultural relevance.
Appendix G Full results
@@
Google Translate Languages and their cousins
As mentioned in the results section, some languages see only very small improvements from finetuning on Smol, and others even see losses. These are mainly either high-resource languages, or close relatives to higher-resource languages. In the full table LABEL:tab:full-results below, The superscript indicates a language supported by Google Translate at the time of these experiments; a superscript like means that this language is closely related to the Google-Translate-supported language xx. We only consider the 108 languages that were present on Google Translate at the time of this work.
| ee | Both | +36.1 | 3.0 | 37.7 | 37.6 | 39.1 | 39.2 | 37.8 | 40.0 | 39.5 | 7.5 | 42.7 | 40.7 |
| kr | Both | +10.8 | 17.3 | 25.6 | 25.9 | 28.1 | 28.8 | 22.7 | 26.3 | 20.3 | 22.2 | 32.6 | 31.0 |
| kg | Both | +9.2 | 34.9 | 46.9 | 36.8 | 44.1 | 43.2 | 43.2 | 47.0 | 37.8 | 29.1 | 50.2 | 3.4 |
| bem | Both | +7.3 | 40.0 | 44.8 | 44.7 | 47.3 | 49.2 | 43.3 | 47.7 | 42.3 | 33.3 | 49.7 | 41.8 |
| dyu | Both | +5.3 | 17.9 | 22.5 | 23.3 | 23.2 | 23.7 | 23.9 | 24.4 | 21.0 | 4.5 | 22.4 | 12.5 |
| din | Both | +4.6 | 20.3 | 23.8 | 22.9 | 24.9 | 25.1 | 23.3 | 25.9 | 21.4 | 1.6 | 25.1 | 26.5 |
| luo | Both | +4.1 | 37.4 | 39.1 | 41.1 | 41.5 | 42.0 | 39.1 | 42.0 | 39.6 | 36.1 | 41.3 | 39.5 |
| fon | Both | +3.6 | 21.3 | 24.3 | 23.9 | 24.9 | 25.3 | 20.4 | 23.7 | 23.8 | 1.9 | 25.9 | 24.2 |
| bm | Both | +3.4 | 30.8 | 28.6 | 35.2 | 34.2 | 34.1 | 34.0 | 36.2 | 33.9 | 9.0 | 35.7 | 32.2 |
| ak | Both | +2.6 | 35.5 | 36.1 | 37.8 | 38.1 | 38.2 | 34.4 | 38.1 | 37.3 | 32.2 | 34.5 | 33.3 |
| ln | Both | +2.5 | 46.8 | 48.1 | 48.4 | 49.3 | 49.3 | 44.6 | 48.3 | 46.5 | 45.2 | 46.4 | 45.7 |
| wo | Both | +1.1 | 30.3 | 30.0 | 30.4 | 31.4 | 31.6 | 31.4 | 32.2 | 30.7 | 29.8 | 36.2 | 30.9 |
| ff | Both | +0.9 | 25.0 | 24.4 | 26.4 | 25.9 | 26.5 | 25.7 | 26.1 | 25.2 | 2.5 | 25.9 | 27.1 |
| om | Both | -0.8 | 40.1 | 38.0 | 39.0 | 39.3 | 39.4 | 39.0 | 40.2 | 41.3 | 38.4 | 41.4 | 39.1 |
| lg | Both | -1.1 | 42.5 | 39.9 | 41.4 | 41.4 | 41.7 | 42.0 | 43.1 | 43.5 | 41.0 | 43.6 | 41.1 |
| ber | Both | -3.2 | 25.3 | 20.6 | 22.9 | 22.1 | 21.9 | 28.5 | 25.2 | 31.1 | 2.8 | 21.0 | 32.4 |
| trp | SmolDoc | +29.7 | 8.4 | 6.5 | 37.8 | 38.1 | 39.1 | 24.7 | 35.8 | 27.2 | 20.3 | 35.9 | - |